# Shortcut Bias Mitigation via Ensemble Diversity Using Diffusion Probabilistic Models

### 23 Nov 2023 v1

@@ Abstract
Spurious correlations in the data, where multiple cues are predictive of the target labels, often lead to a phenomenon known as simplicity bias, where a model relies on erroneous, easy-to-learn cues while ignoring reliable ones. In this work, we propose an ensemble diversification framework exploiting Diffusion Probabilistic Models (DPMs) for shortcut bias mitigation. We show that at particular training intervals, DPMs can generate images with novel feature combinations, even when trained on images displaying correlated input features. We leverage this crucial property to generate synthetic counterfactuals to increase model diversity via ensemble disagreement. We show that DPM-guided diversification is sufficient to remove dependence on primary shortcut cues, without a need for additional supervised signals. We further empirically quantify its efficacy on several diversification objectives, and finally show improved generalization and diversification performance on par with prior work that relies on auxiliary data collection.
@@ 1 Introduction
Addressing the simplicity bias in machine learning models has been a focus of extensive research. Numerous studies have aimed to encourage models to use a broader and more diversified set of predictive cues, especially when dealing with training data that lacks explicit shortcut cue labels. A variety of these methods have been input-centric, designed to drive models to focus on different areas of the input space [26, 14], while others have focused on diversification strategies that rely on auxiliary data for disagreement [16, 9]. The latter approach, in particular, has been instrumental in developing functionally diverse models that exhibit robustness to shortcut biases, albeit with the limitation of requiring access to expensive auxiliary data.
[…]
@@ 2.1 Overcoming the Simplicity Bias
To overcome the simplicity bias, various methods have strived to find ways to drive models to use a diversified set of predictive signals. Copious literature has explored methodologies to avoid or mitigate shortcut cue learning when labels for the shortcut cues were present in the training data [10, 7, 28, 6, 18]. As it is not always easy to predict shortcut signals, it is generally hard or impossible to have appropriate labels associated with the shortcut cues.
A common solution has historically been to train an ensemble of models, where each predictor is encouraged to use different cues for inference, thus increasing the chance for some to avert their attention from simple cues. To ensure model diversity, many methods have been proposed. For example, prior work has shown how a set of models can be trained while ensuring mutual orthogonality of the input gradients, effectively driving models to attend to different locations of the input space for prediction [17, 25, 26, 14]. These input-centric methods, however, may be at a disadvantage in cases where different features must attend to the same area of the input space.
Instead, a more direct approach has been via the direct diversification of the models’ outputs within the ensemble. This approach hinges on the availability of –unlabelled– auxiliary, and effectively Out-Of-Distribution (ood), data that is, at least in part, free of the same shortcuts as the original training data. Through a diversification objective, then, the models are made to disagree on these the ood samples, effectively fitting functions with different extrapolation behaviors [9, 16, 22]. Even in this case, the dependency on auxiliary ood data poses some limitations, as this is often not readily accessible, and can be costly to procure.
@@ 2.2 DPMs Modelling and Generalization
In recent years, Diffusion Probabilistic Models have emerged as a transformative generative tool. Numerous studies have underscored their prowess in generating synthetic images, which can then be harnessed to enrich datasets and bolster classification performance [20, 2, 33]. In several cases, DPMs have been shown to transcend the surface-level statistics of the data, making them invaluable in understanding data distributions and features [4, 33, 30]. Moreover, recent studies have indicated the ability of DPMs to achieve feature disentanglement via denoising reconstructions [8, 30, 15]. In particular, work in [29] has shown how text-guided generative models can represent disentangled concepts, and how through algebraic manipulation of their latent representations it is possible to compositionally generalize to novel and unlikely image feature combinations. In our work, we test the edge case of DPMs trained with data exhibiting correlated input features, and when suitably trained, we find them capable of generating samples with novel feature combinations.
@@ 3.2 Datasets
As the experimental grounds for our study, we leverage two representative datasets, a color-augmented version of DSprites [12] and UTKFace [35], previously shown to contain features leading to strong cue learning biases [21].
DSprites: DSprites includes a comprehensive set of symbolic objects generated with variations in five latent variables: shape, scale, orientation, and X-Y position. We augment this dataset with a color dimension and remove redundant samples (e.g. due to rotations), resulting in $2,862,824$ distinct images, which we refer to as ColorDSprites. Within the context of WCST-ML, ColorDSprites permits the examination of shortcut biases in a highly controlled setup.
@@ 3.3 DPMs and Efficient Sampling
To facilitate efficient sampling, we employ Denoising Diffusion Implicit Models (DDIM) [24], a first-order ODE solver for DPMs [19, 11], utilizing a predictor-corrector scheme to minimize the number of sampling step. In particular, we set $\sigma_{t}=0$ in the predictor-corrector scheme to enhance the efficiency of the sampling process.
In our setup, we consider DPMs trained at different fidelity levels, postulating that, generally, DPMs trained for longer durations should more closely fit the data distribution at hand. Therefore, in our experiments, we use the ‘number of diffusion training epochs’ as a proxy for the DPM fidelity on the modeled distribution.
@@ 3.4 Ensemble Training and Diversification
$\mathcal{L}^{\text{div}}_{\text{reg}}=\sum_{i\neq j}$ $D_{KL}(p(f_{i},f_{j})||p(f_{i})\otimes p(f_{j}))$ (6)
$+\sum_{i}D_{KL}(p(f_{i})||p(y))$
@@
3.5 Experiment Design
For each of the ensemble experiments, we train a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. We perform ensemble training separately with all diversification objectives considered, and for each objective, we perform ablation studies applying the diversification objectives to the data generated by the DPMs at different fidelity levels. For comparison, we also consider the ensemble diversification by using real ood data, randomly sampled from the off-diagonal sets, as well as a baseline without a diversification objective.
4 Results
@@
4.1 Diffusion Disentanglement & OOD sampling
DPMs Exhibit Generalization Capabilities Under Correlation:
We test the ability of DPMs to transcend surface-level statistics of the data and generate samples that break the shortcut signals at training time, even when trained on correlated data. We train the DPMs on the correlated WCST-ML diagonal sample sets for both ColorDSprites and UTKFace. Fig. 2 displays the training samples for both datasets (left halves), as well as samples from the DPMs (right halves) trained for respectively 25 and 800 epochs. In the figure we observe how sampling from the trained DPMs generates previously unseen feature combinations, despite the correlated coupling of features during training (e.g. $\langle green/white,heart\rangle$, $\langle green,square\rangle$ or $\langle child,female\rangle$), suggesting a potential for innate disentangling capabilities in diffusion latent spaces. Leveraging this powerful capability is crucial in the generation of samples for disagreement, as it allows models to break from the shortcuts in the training distribution.
Early Stopping to Capture Diffusion OOD Sampling Capabilities:
To understand under which conditions sampling from DPMs can lead to the generation of samples displaying unseen feature combinations, we examine the fraction of ood samples generated by DPMs at different fidelities (§3.5). We consider ColorDSprites, given its synthetic nature, and the minimal overlap of different feature sets, making it particularly useful for this analysis. To measure the fraction of ood samples, we train a near-perfect oracle on the full ColorDSprites dataset, trained to predict the WCST-ML partitioned labels under all features. We can then consider in-distribution (id) those samples showcase fully correlated feature levels (close to the diagonal samples), and out-of-distribution (ood) those samples classified to belong to the off-diagonal WCST-ML set. Fig. 3 shows the fraction of OOD samples generated by the ColorDSprites DPM at varying levels of training fidelities. We identify at least three qualitative different intervals. An initial burn-in interval, characterized by a high frequency of ood generated samples, but which fails to capture the manifold of the data under inspection; an originative interval, where we observe a reduced number of ood samples, in favor for a generative distribution more aligned with the data to represent; and an exact interval, where the DPM’s ability to almost perfectly represent the data comes at the cost of novel emergent feature mixtures. In the context of ensemble diversification and bias mitigation, the originative interval is posed to provide the necessary information to break the simplicity shortcuts, while providing effective disagreement signals with respect to the visual cues available.
4.2 Diffusion-guided Ensemble Diversity
We trained a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. The ablation studies examined the training of the ensemble with each diversification objective in turn, as well as a baseline with no imposed diversification. For both baseline and diversification experiments, we favor different training dynamics by using separate vanilla Adam optimizers for each model. We train the ensemble with a cross-entropy loss on the correlated diagonal training data, as well as an additional, objective-specific, diversification loss, computed on a separate diversification set. In our experiments, we will refer to ‘ood’ when the data for disagreement is comprised of left-out, feature-uncorrelated, samples from the original dataset; to ‘diffusion’ when the samples are generated by a DPM; and to ‘baseline’ when no diversification objective is included. We tune $\gamma$ by grid-search to provide comparable experiments under different objectives (see Suppl. §6.1)
@@ 4.2 Diffusion-guided Ensemble Diversity
DPM Fidelity Significantly Impacts Diversification:
Diffusion-guided Diversity Leads to Comparably Diverse Ensembles:
Given the difficulty and cost of collecting auxiliary ood data for diversification, we wish to compare the level of diversification achieved with DPM counterfactual, as opposed to using real ood data. Based on our previous results in Fig. 4 (see also Suppl. Fig. S2), which do not require access to auxiliary data, we choose early-stopping DPM training at 100 and 800 epochs for ColorDSprites and UTKFace respectively, providing samples especially useful for diversification. In Fig. 5 we report the objective-wise normalized diversity achieved in each scenario by the ensemble. We find that diffusion counterfactuals can lead to almost comparable diversification performance with respect to real ood samples. In the figure, the diffusion-led diversity is almost always within 5% from the metrics achieved when using pure ood samples, both typically over 50% higher than the baseline.
@@
Ensemble Diversification Breaks Simplicity Biases:
By WCST-ML, we can test each model’s bias to a cue by testing the model’s output on purposefully designed test sets [21]. We test the quality of the diversification obtained in Table 1, and ascertained the fraction of models attending to specific cues when trained on out-of-distribution and diffusion-generated samples respectively. A model was deemed to be attending to a cue if its validation accuracy on the cue-specific ood WCST-ML dataset was highest relative to all other features. Our baseline findings mirrored the observations in [21]. Specifically, each model in the baseline ensemble predominantly attended the color cue in ColorDSprites and the ethnicity cue in UTKFace, showcasing strong preferential cue bias while achieving near-perfect classification on diagonal –id– validation data. Notably, upon introducing the diversification objectives, we observed a perceptible shift in the models’ behavior, some of which averted their focus from the primary, easy-to-learn cues, turning instead to other latent cues present within the data. Among the objectives considered, $kl$, $l_{1}$, and $l_{2}$ exhibited the highest cue diversity, catalyzing the ensemble to distribute attention across multiple cues. However, this also came at the expense of a marked drop in the average ensemble performance. Conversely, the div and cross objectives yielded milder diversification, focusing primarily on the next readily discernible cues, specifically scale for ColorDSprites and gender for UTKFace, all the while maintaining a generally higher ensemble validation performance. As confirmed in Fig. 5, the diversification achieved by Diffusion counterfactuals in Table 1 is largely comparable with ood data on ColorDSprites, with respectively up to $11\%$ and $14\%$ of the models, averting their attention from the main ‘color’ cue in ColorDSprites, and respectively up to $23\%$ and $36.8\%$ of the models averting their attention to the ethnicity cue in UTKFace.
Increased Ensemble Disagreement Negatively Correlates with Ensemble id Performance
Achieving ensemble diversity by disagreement on unlabelled ood samples has previously been shown to negatively impact id ensemble performance [16] (suppl. Fig. S1), while improving ood performance on downstream tasks associated with non-shortcut cues. Generally, ensemble model selection has been an effective method to prune models that misalign with the original classification objective [9]. In this section, we wish to understand this relationship under the scope of Diffusion-guided diversification. In Fig. 6 we observe the ensemble average accuracy $\delta$ as a function of the diversity $\delta$, spanning all the diversification objectives. We highlight a few important observations. First, increased ensemble diversity via disagreement negatively correlates with average ensemble accuracy, without additional model selection mechanisms. Second, real ood data achieves the highest diversity gain, comparable to the diffusion samples at the originative stage (§4.1), $\approx[10,100]$ for ColorDSprites, and $\approx[450,800]$ for UTKFace. Finally, the diversification objectives show comparable trends, with the $div$ objective displaying marginally better diversification/accuracy performance than the others on both tasks.

### 05 Mar 2024 v2

# Mitigating Biases with Diverse Ensembles and Diffusion Models

@@ Abstract
Spurious correlations in the data, where multiple cues are predictive of the target labels, often lead to a phenomenon known as shortcut bias, where a model relies on erroneous, easy-to-learn cues while ignoring reliable ones. In this work, we propose an ensemble diversification framework exploiting Diffusion Probabilistic Models (DPMs) for shortcut bias mitigation. We show that at particular training intervals, DPMs can generate images with novel feature combinations, even when trained on samples displaying correlated input features. We leverage this crucial property to generate synthetic counterfactuals to increase model diversity via ensemble disagreement. We show that DPM-guided diversification is sufficient to remove dependence on primary shortcut cues, without a need for additional supervised signals. We further empirically quantify its efficacy on several diversification objectives, and finally show improved generalization and diversification performance on par with prior work that relies on auxiliary data collection.
Keywords:
Bias Mitigation Ensemble Diversity Diffusion Probabilistic Models
@@ 1 Introduction
Addressing the simplicity bias in machine learning models has been a focus of extensive research. Numerous studies have aimed to encourage models to use a broader and more diversified set of predictive cues, especially when dealing with training data that lacks explicit shortcut cue labels. A variety of these methods have been input-centric, designed to drive models to focus on different areas of the input space [33, 21], while others have focused on diversification strategies that rely on auxiliary data for disagreement [23, 13]. The latter approach, in particular, has been instrumental in developing functionally diverse models that exhibit robustness to shortcut biases, albeit limited by their required access to expensive auxiliary data.
@@ 2.1 Overcoming the Simplicity Bias
To overcome the simplicity bias, copious literature has explored methodologies to avoid or mitigate shortcut cue learning when labels for the shortcut cues were present in the training data [14, 10, 35, 8, 25, 12]. The access to shortcut signals is however a critical limitation, as these are generally hard or impossible to obtain. Data augmentation methodologies have proven useful in the generation of bias-conflicting or bias-free samples [9, 15, 6], or to augment underrepresented subgroups therein [36, 20], leading to debiased predictions. Although this is an important research direction, the missed reliance on shortcut cues for prediction may not always lead to improved generalization, notably when the downstream task is aligned with the shortcut cue.
A different approach to this problem has instead been to enforce the use of a diverse set of signals for prediction. The use of ensembles has been one such method, where models’ diversity would lead to bias mitigation and improved generalization. Different weight initialization and architectures have previously been shown ineffective in the presence of strong shortcut biases [28]. Methods ensuring mutual orthogonality of the input gradients have proven more effective, driving models to attend to different locations of the input space for prediction [24, 32, 33, 21]. These input-centric methods, however, may be at a disadvantage in cases where different features must attend to the same area of the input space. Instead, a different approach has been via the diversification of direct ensemble model predictions. This approach hinges on the availability of –unlabelled– out-of-distribution (ood) auxiliary samples that are, at least in part, free of the same shortcuts as the original training data. Through a diversification objective, then, the models are made to disagree on these the ood samples, effectively fitting functions with different extrapolation behaviors [13, 23, 29, 16]. Even in this case, the auxiliary ood data dependency poses limitations, as this is often not readily accessible, and can be costly to procure.
@@ 2.2 DPMs Modelling and Generalization
In recent years, Diffusion Probabilistic Models have emerged as a transformative generative tool. Numerous studies have underscored their prowess in generating synthetic images, which can then be harnessed to enrich datasets and bolster classification performance [27, 2, 41]. In several cases, DPMs have been shown to transcend the surface-level statistics of the data, making them invaluable in understanding data distributions and features [4, 41, 38]. Moreover, recent studies have indicated the ability of DPMs to achieve feature disentanglement via denoising reconstructions [11, 38, 22]. In particular, work in [37] has shown how text-guided generative models can represent disentangled concepts, and how through algebraic manipulation of their latent representations it is possible to compositionally generalize to novel, and unlikely, image feature combinations. Additional work has also shown strong inductive Diffusion biases during learning, which may explain some of these phenomena [7]. In our work, we test the edge case of DPMs trained with data exhibiting correlated input features. We find that when suitably trained, DPMs can still generate samples with novel feature combinations and that these can be leveraged for ensemble diversity.
@@ 3.2 Datasets
As the experimental grounds for our study, we leverage two representative datasets: a color-augmented version of DSprites [18], and UTKFace [43]. The choice of the datasets in our work was due to several factors; most prominently, the disentangled nature of the features in DSPrites – providing a particularly useful analysis with a controlled overlap of bias tendencies– and a more complex UTKFace, previously known to present strong ethnical bias in WCST-ML training, with concerning societal implications [28].
DSprites: DSprites includes a comprehensive set of symbolic objects generated with variations in five latent variables: shape, scale, orientation, and X-Y position. We augment this dataset with a color dimension and remove redundant samples (e.g. due to rotations), resulting in $2,862,824$ distinct images, which we refer to as ColorDSprites. ColorDSprites permits the examination of shortcut biases in a highly controlled setup.
@@
We train our models on the correlated WCST-ML diagonal sample sets for both ColorDSprites and UTKFace. For ColorDSprites, we consider $K_{DS}=4$, features $\{color,orientation,scale,shape\}$, and $L=3$ as constrained by the number of shapes in the dataset. Within UTKFace we consider $K_{UTK}=3$, features $\{ethnicity,gender,age\}$, and $L=2$ as constrained by the binary classification on gender. See §S1 for additional implementation details.
@@ 3.3 DPMs and Efficient Sampling
To facilitate efficient sampling, we employ Denoising Diffusion Implicit Models (DDIM) [31], a first-order ODE solver for DPMs [26, 17], utilizing a predictor-corrector scheme to minimize the number of sampling step. In particular, we set $\sigma_{t}=0$, to enhance the efficiency of the sampling process.
In our setup, we consider DPMs trained at different fidelity levels, postulating that, generally, increased DPM training will more closely fit the data distribution at hand. Therefore, in our experiments, we use the ‘number of diffusion training epochs’ as a proxy for the DPM fidelity on the modeled distribution.
@@ 3.4 Ensemble Training and Diversification
$\mathcal{L}^{\text{div}}_{\text{reg}}=\sum_{i\neq j}D_{KL}(p(f_{i},f_{j})||p(f_{i})\otimes p(f_{j}))+\sum_{i}D_{KL}(p(f_{i})||p(y))$ (6)
@@
3.5 Experiment Design
For each of the ensemble experiments, we train a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. We perform ensemble training separately with all diversification objectives considered, and for each objective, we perform ablation studies applying the diversification objectives to the data generated by the DPMs at different fidelity levels. For comparison, we also consider ensemble diversification with real ood data, randomly sampled from the off-diagonal sets, as well as a baseline without a diversification objective.
In summary, the full analysis in this work involved the training of over 30 diffusion models (UNet $\approx 35$million parameters), and $\approx 16,000$ ResNet-18 classifiers (each $\approx 11$million parameters), sectioned into 160 ensembles (one for each diffusion-objective pair and baselines), and totaling the training of over 177 billion learnable parameters.
4 Results
4.1 Diffusion Disentanglement & ood sampling
DPMs Exhibit Generalization Capabilities Under Correlation.
We test the ability of DPMs to transcend surface-level statistics of the data and generate samples that break the shortcut signals at training time, even when trained on correlated data. Fig. 2 displays the training samples for both datasets (left halves), as well as samples from the DPMs (right halves) trained for respectively 25 and 800 epochs. In the figure we observe how sampling from the trained DPMs generates previously unseen feature combinations, despite the correlated coupling of features during training (e.g. $\langle green/white,heart\rangle$, $\langle green,square\rangle$ or $\langle child,female\rangle$), suggesting a potential for innate disentangling capabilities in diffusion latent spaces. Leveraging this powerful capability is crucial in the generation of samples for disagreement, as it allows models to break from the shortcuts in the training distribution.
Early Stopping to Capture Diffusion ood Sampling Capabilities.
To understand under which conditions sampling from DPMs can lead to the generation of samples displaying unseen feature combinations, we examine the fraction of ood samples generated by DPMs at different fidelities (§3.5). The ColorDSprites is particularly useful for this analysis, given the disentangled nature of its features. To measure the fraction of ood samples, we train a near-perfect oracle on the full ColorDSprites dataset, trained to predict the WCST-ML partitioned labels under all features. We can then consider in-distribution (id) those samples showcase fully correlated feature levels (close to the diagonal samples), and out-of-distribution (ood) those samples classified to belong to the off-diagonal WCST-ML set. Fig. 3 shows the fraction of ood samples generated by the ColorDSprites DPM at varying levels of training fidelities. We identify at least three qualitative different intervals. An initial burn-in interval, characterized by a high frequency of ood generated samples, but which fails to capture the manifold of the data under inspection; an originative interval, where we observe a reduced number of ood samples, in favor for a generative distribution more aligned with the data to represent; and an exact interval, where the DPM’s ability to almost perfectly represent the data comes at the cost of novel emergent feature mixtures. In the context of ensemble diversification and bias mitigation, the originative interval is posed to provide the necessary information to break the simplicity shortcuts, while providing effective disagreement signals with respect to the visual cues available. We refer the reader to Suppl. §S2 for further insights into DPM-early stopping with ensemble diversification criteria.
@@ 4.1 Diffusion Disentanglement & ood sampling
4.2 Diffusion-guided Ensemble Diversity
We trained a diverse ensemble comprising 100 ResNet-18 models on both ColorDSprites and UTKFace. The ablation studies examined the training of the ensemble with each diversification objective in turn, as well as a baseline with no imposed diversification. For both baseline and diversification experiments, we favor different training dynamics by using separate vanilla Adam optimizers for each model. We train the ensemble with a cross-entropy loss on the correlated diagonal training data, as well as an additional, objective-specific, diversification loss, computed on a separate diversification set. In our experiments, we will refer to ‘ood’ when the data for disagreement is comprised of left-out, feature-uncorrelated, samples from the original dataset; to ‘diffusion’ when the samples are generated by a DPM; and to ‘baseline’ when no diversification objective is included. We tune $\gamma$ by grid-search to provide comparable experiments under different objectives (see Suppl. §S2.1)
DPM Fidelity Significantly Impacts Diversification.
Diffusion-guided Diversity Leads to Comparably Diverse Ensembles.
@@ 4.2 Diffusion-guided Ensemble Diversity
Given the difficulty and cost of collecting auxiliary ood data for diversification, we wish to compare the level of diversification achieved with DPM counterfactual, as opposed to using real ood data. Based on our previous results in Fig. 4 (also Suppl. Fig. S4), we choose early-stopping DPM training at 100 and 800 epochs for ColorDSprites and UTKFace respectively, providing samples especially useful for diversification. In Fig. 5 we report the objective-wise normalized diversity achieved in each scenario by the ensemble. We find that diffusion counterfactuals can lead to comparable diversification performance with respect to real ood samples. In the figure, the diffusion-led diversity is almost always within 5% from the metrics achieved when using pure ood samples, both typically over 50% higher than the baseline.
Ensemble Diversification Breaks Simplicity Biases.
@@
By WCST-ML, we can test each model’s bias to a cue by testing the model’s output on purposefully designed test sets [28]. We test the quality of the diversification obtained in Table 1, and ascertained the fraction of models attending to specific cues when trained on out-of-distribution and diffusion-generated samples respectively. A model was deemed to be attending to a cue if its validation accuracy on the cue-specific ood WCST-ML dataset was highest relative to all other features. Our baseline findings mirrored the observations in [28]. Specifically, each model in the baseline ensemble predominantly attended the color cue in ColorDSprites and the ethnicity cue in UTKFace, showcasing strong preferential cue bias while achieving near-perfect classification on diagonal –id– validation data. Notably, upon introducing the diversification objectives, we observed a perceptible shift in the models’ behavior, some of which averted their focus from the primary, easy-to-learn cues, turning instead to other latent cues present within the data. Among the objectives considered, $kl$, $l_{1}$, and $l_{2}$ exhibited the highest cue diversity, catalyzing the ensemble to distribute attention across multiple cues. However, this also came at the expense of a marked drop in the average ensemble performance. Conversely, the div and cross objectives yielded milder diversification, focusing primarily on the next readily discernible cues, specifically scale for ColorDSprites and gender for UTKFace, all the while maintaining a generally higher ensemble validation performance. As confirmed in Fig. 5, the diversification achieved by Diffusion counterfactuals in Table 1 is largely comparable with ood data on ColorDSprites, with respectively up to $11\%$ and $14\%$ of the models averting their attention from the main ‘color’ cue in ColorDSprites, and up to $23\%$ and $36.8\%$ of the models averting their attention to the ethnicity cue in UTKFace.
Increased Ensemble Disagreement Negatively Correlates with Ensemble id Performance.
@@
Achieving ensemble diversity by disagreement on unlabelled ood samples has previously been shown to negatively impact id ensemble performance [23] (suppl. Fig. S1), while improving ood performance on downstream tasks associated with non-shortcut cues. Generally, ensemble model selection has been an effective method to prune models that misalign with the original classification objective [13]. In this section, we wish to understand this relationship under the scope of Diffusion-guided diversification. In Fig. 6 we observe the ensemble average accuracy $\delta$ as a function of the diversity $\delta$, spanning all the diversification objectives. We highlight a few important observations. First, increased ensemble diversity via disagreement negatively correlates with average ensemble accuracy, without additional model selection mechanisms. Second, real ood data achieves the highest diversity gain, comparable to the diffusion samples at the originative stage (§4.1), $\approx[10,100]$ for ColorDSprites, and $\approx[450,800]$ for UTKFace. Lastly, the diversification objectives show comparable trends, with the $div$ objective displaying marginally better diversification/accuracy performance than the others on both tasks.